# S-Duality for Non-Abelian Monopoles

### v1

@@ 5.1 $(H^{\vee})^{s}$-weight decomposition of harmonic forms
    $\left\{\tilde{\omega}^{j-i}\wedge\tilde{\Omega}^{i-1}\;\middle|\;1\leq i\leq j\leq N+1\right\}$ (5.75)
    $=$ $\left\{\tilde{\omega}^{j-1}\;\middle|\;1\leq j\leq N+1\right\}\cup\left\{\tilde{\Omega}\wedge(\tilde{\omega}^{j-i-1}\wedge\tilde{\Omega}^{i-1})\;\middle|\;1\leq i<j\leq N+1\right\},$
    where $\tilde{\omega}^{j-1}$ and $\tilde{\omega}^{j-i-1}\wedge\,\tilde{\Omega}^{i-1}$ arise from $C(k_{1})$ and $C(k_{2})$, respectively. Here $\text{dim}\tilde{\Omega}=4=\text{dim}\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0},k_{2})$ as in (5.11). The weight assignment is
    $\text{wt}(f_{ij})=\begin{cases}\text{wt}_{jj},&i=1,\;1\leq j\leq N+1,\\[4.0pt] \text{wt}_{i-1\;j},&2\leq i\leq j\leq N+1.\end{cases}$ (5.76)
    The simple root generators of $SU(N+1)$ in this basis are constructed as
    $E_{\bm{\alpha}_{k}}$ $=\sum_{a=1}^{k-1}E_{(N+3-k)\;(N+2-a)\,,\,(N+2-k)\;(N+2-a)}+\sqrt{2}\,E_{1\;(N+2-k)\,,\,(N+2-k)\;(N+2-k)}$ (5.77)
    $\quad+\sum_{b=k+2}^{N+1}E_{(N+3-b)\;(N+2-k)\,,\,(N+3-b)\;(N+1-k)}+\sqrt{2}\,E_{(N+2-k)\;(N+2-k)\,,\,1\;(N+1-k)},$
    where $E_{ij,pq}$ are $E_{m,n}$ defined in (4.61), with the single-index basis $f_{m}$ replaced by the double-index basis $f_{ij}$.
@@ 5.2 Non-abelian electric and magnetic gauge transformations
In $\mathcal{N}=4$ SYM with the symmetry breaking $G\rightarrow H^{s}\times U(1)$, $H^{s}$ acts isometrically on the relative moduli space $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0})$, with generators given by Killing vector fields. The Hilbert space of states on $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0})$ also admits an action of the dual group $(H^{\vee})^{s}$. While $H^{s}$ acts geometrically by metric-preserving diffeomorphisms, $(H^{\vee})^{s}$ acts algebraically by shifting weights within the representation space. On differential forms, this algebraic action is realized by wedge ($O[f]$) and contraction ($O^{\dagger}[f]$) operations.
@@
In what follows, we construct the explicit $H^{s}\times(H^{\vee})^{s}$-representation on the Hilbert space of $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$. In this case all harmonic forms are parallel, so a primitive (Lefschetz) decomposition is available.
@@
where each multiplet $\{\beta_{\alpha}^{(j_{k})}\}_{k=1}^{r_{\alpha}^{(j)}}$ consists of forms of the same degree. Then $\mathcal{H}_{\alpha}$ is spanned by
$\Big\{\,L_{1}^{i_{1}}\cdots L_{s}^{i_{s}}\,\beta_{\alpha}^{(1_{k})}\;\Big|\;i_{m}\in\mathbb{Z}_{\geq 0},1\leq k\leq r^{(1)}_{\alpha}\,\Big\}\cup\Big\{\,L_{1}^{i_{1}}\cdots L_{s}^{i_{s}}\,\beta_{\alpha}^{(2_{k})}\;\Big|\;i_{m}\in\mathbb{Z}_{\geq 0},1\leq k\leq r^{(2)}_{\alpha}\,\Big\}$ (5.86)
$\cup\cdots\cup\Big\{\,L_{1}^{i_{1}}\cdots L_{s}^{i_{s}}\,\beta_{\alpha}^{(n_{k})}\;\Big|\;i_{m}\in\mathbb{Z}_{\geq 0},1\leq k\leq r^{(n)}_{\alpha}\,\Big\}$.
@@ 6 Discussion
In a unit-charge sector, where $m$ has a single nonzero component equal to $1$, the dynamics reduces to a free particle, with no interactions. For maximal symmetry breaking $G\to U(1)^{r}$, $\mathcal{M}(m,\Phi_{0})=\mathbb{R}^{3}\times\mathbb{S}^{1}$, so semiclassical quantization yields a particle with no internal structure. When $t<r$, the relative moduli space $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0})$ is nontrivial, and the magnetic particle carries internal degrees of freedom, on which the magnetic gauge symmetry acts.
The simplest case is $SU(N+1)\to U(N)$, where the relative moduli space is the magnetic orbit $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=C(k_{1})\cong\mathbb{CP}^{N-1}$, with position eigenstates $|G_{0}\rangle$. Each $|G_{0}\rangle$ represents a classical configuration with magnetic charge $G_{0}$. For BPS particles, the internal states are $N$ harmonic wavefunctions on $C(k_{1})$ denoted by $\{{|{k}\rangle}\}_{k=1}^{N}$, which form an $SU(N)$ isospin multiplet. The electric $SU(N)$ action moves $|{G_{0}}\rangle$} along the orbit, leaving $\{{|{k}\rangle}\}_{k=1}^{N}$ fixed. By contrast, the magnetic $SU(N)$ action rotates the BPS isospin multiplet and leaves $G_{0}$ fixed: $|G_{0}\rangle|\Omega\rangle\rightarrow|G_{0}\rangle\exp\{i\theta(G_{0},\hat{\lambda}^{\dagger},\hat{\lambda})\}|\Omega\rangle$.
@@ Appendix C Effective Lagrangian via the collective-coordinate expansion
To evaluate the $\bar{\Psi}\,\gamma^{I},\Psi]$ contribution, we return to the eight real spinors $\varepsilon^{\mathcal{A}}_{0}$ $(\mathcal{A}=1,\ldots,8)$ defined in Section 4, and set $\mathcal{A}=1,5$ at the end to recover $\varepsilon^{\alpha}_{0}$. Let

### v2

@@ A Schubert basis of harmonic forms is
    $\left\{(f_{ij})_{\mathrm{harm}}\mid 1\leq i\leq j\leq N+1\right\}$ (5.75)
    $=$ $\left\{(f_{1j})_{\mathrm{harm}}\mid 1\leq j\leq N+1\right\}\cup\left\{(\tilde{\Omega}\wedge f^{\prime}_{ij})_{\mathrm{harm}}\mid 1\leq i<j\leq N+1\right\},$
    where the restrictions $\{[f_{1j}]|_{C(k_{1})}\}$ form a basis of $H^{\ast}(C(k_{1}))$, while $\{[\tilde{\Omega}\wedge f^{\prime}_{ij}]\}$ is the Thom-shifted image of the Schubert basis of $H^{\ast}(C(k_{2}))$. Here $\deg(\tilde{\Omega})=4=\text{dim}\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0},k_{2})$ as in (5.11). The weight assignment is
    $\text{wt}((f_{ij})_{\text{harm}})=\begin{cases}\text{wt}_{jj},&i=1,\;1\leq j\leq N+1,\\[4.0pt] \text{wt}_{i-1\;j},&2\leq i\leq j\leq N+1.\end{cases}$ (5.76)
@@ 5.2 Non-abelian electric and magnetic gauge transformations
In $\mathcal{N}=4$ SYM with the symmetry breaking $G\rightarrow H^{s}\times U(1)^{t}$, $H^{s}$ acts isometrically on the relative moduli space $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0})$, with generators given by Killing vector fields. The Hilbert space of states on $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0})$ also admits an action of the dual group $(H^{\vee})^{s}$. While $H^{s}$ acts geometrically by metric-preserving diffeomorphisms, $(H^{\vee})^{s}$ acts algebraically by shifting weights within the representation space. On differential forms, this algebraic action is realized by wedge ($O[f]$) and contraction ($O^{\dagger}[f]$) operations.
@@
In what follows, we construct the explicit $H^{s}\times(H^{\vee})^{s}$-representation on the Hilbert space of $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=C(k_{1})$. In this case all harmonic forms are parallel, so a primitive (Lefschetz) decomposition is available.
@@
where the forms in each $\{\beta_{\alpha}^{(j_{k})}\}_{k=1}^{r_{\alpha}^{(j)}}$ have the same degree. Then $\mathcal{H}_{\alpha}$ is spanned by
$\bigcup^{n}_{j=1}\Big\{\,L_{1}^{i_{1}}\cdots L_{s}^{i_{s}}\,\beta_{\alpha}^{(j_{k})}\;\Big|\;i_{m}\in\mathbb{Z}_{\geq 0},\,1\leq k\leq r^{(j)}_{\alpha}\,\Big\}$. (5.85)
@@ 6 Discussion
In a unit-charge sector, where $m$ has a single nonzero component equal to $1$, the dynamics reduces to a free particle, with no interactions. For maximal symmetry breaking $G\to U(1)^{r}$, $\mathcal{M}(m,\Phi_{0})=\mathbb{R}^{3}\times\mathbb{S}^{1}$, so semiclassical quantization yields a particle with no internal structure. When $t<r$, the relative moduli space $\mathcal{M}_{\mathrm{rel}}(m,\Phi_{0})$ can be nontrivial, and the magnetic particle may carry internal degrees of freedom, on which the magnetic gauge symmetry acts.
The simplest case is $SU(N+1)\to U(N)$, where the relative moduli space is the magnetic charge orbit $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=C(k_{1})\cong\mathbb{CP}^{N-1}$, with position eigenstates $|G_{0}\rangle$. Each $|G_{0}\rangle$ represents a field configuration with magnetic charge $G_{0}$. For BPS particles, the internal states are $N$ harmonic wavefunctions on $C(k_{1})$ denoted by $\{{|{k}\rangle}\}_{k=1}^{N}$, which form an $SU(N)$ isospin multiplet. The electric $SU(N)$ action moves $|{G_{0}}\rangle$} along the orbit, leaving $\{{|{k}\rangle}\}_{k=1}^{N}$ fixed. By contrast, the magnetic $SU(N)$ action rotates the BPS isospin multiplet and leaves $G_{0}$ fixed: $|G_{0}\rangle|\Omega\rangle\rightarrow|G_{0}\rangle\exp\{i\theta(G_{0},\hat{\lambda}^{\dagger},\hat{\lambda})\}|\Omega\rangle$.
@@ Appendix C Effective Lagrangian via the collective-coordinate expansion
To evaluate the $\bar{\Psi}\,\gamma^{I},\Psi]$ contribution, we return to the eight real spinors $\varepsilon^{\mathcal{A}}_{0}$ $(\mathcal{A}=1,\ldots,8)$ in (4.19), and set $\mathcal{A}=1,5$ at the end to recover $\varepsilon^{\alpha}_{0}$. Let